At diseases of a pancreas cabbage juice is used, pH of it is 5,44. To find [H ] and [OH] (mole/1).
1
Expert's answer
2020-09-17T07:16:22-0400

We know pH= -log [H+]

So, here pH = 5.44

So, -log [H+] = 5.44

Or, [H+] = 10^-5.44

= 3.63×10^-6 M

We also know that pH + pOH = 14

So, pOH = 14-pH

Or, pOH= 14-5.44

= 8.56

Now, -log[OH-] = pOH

OR,-log[OH-] = 8.56

Or, [OH-] = 10^-8.56

= 2.75×10^-9 M